OpenStudy (anonymous):

A population of 80,000 toads is expected to shrink at a rate of 9.2% per year. What will the toad population be in 20 years? 3632 7360 11,609 72,640

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (anonymous):

Divide the population by the percent 80,000-9.2%=72640 Then divide your answer by 20 years'.

OpenStudy (ranga):

\(\Large P = 80,000(1 - 0.092)^t \) \(\Large P = 80,000(0.908)^t \) Put t = 20 in the above formula.

OpenStudy (anonymous):

Thank you guys so much!

OpenStudy (anonymous):

Anytime :)

OpenStudy (anonymous):

Okay, so I got the answer 11, 609. Is that right? If not, could you possibly tell me what Im doing wrong so I can understand how to do it correctly.

OpenStudy (ranga):

You have the correct answer.

OpenStudy (anonymous):

Okay, thank you ranga!

OpenStudy (ranga):

You are welcome.

OpenStudy (anonymous):

I'm having a little trouble with this question. In 2010, a city’s population was 1,405,233 and it was decreasing at a rate of 1.1%. At this rate, when will the city’s population fall below 1,200,000? in 2024 in 2027 in 2036 in 2049

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Similar formula as before but this time you have to solve for t. \(\Large P = 1,405,233(1 - 0.011)^t\) \(\Large P = 1,405,233(0.989)^t\) \(\Large 1,200,000 = 1,405,233(0.989)^t\) Solve for 't'.

OpenStudy (ranga):

\(\Large \frac{1,200,000}{1,405,233} = 0.989^t \)

OpenStudy (anonymous):

Would the answer be in 2027?

OpenStudy (ranga):

I am getting a different answer.

OpenStudy (anonymous):

Okay, let me retry this. I am thinking I made a mistake on my calculator that I am not catching.

OpenStudy (ranga):

\( \Large 0.8539509 = 0.989^t \) \( \Large \log(0.8539509) = \log(0.989^t) = t\log(0.989) \) \( \Large t = \frac{\log(0.8539509)}{\log(0.989)} = ? \)

OpenStudy (anonymous):

Okay, so would the answer be in 2036?

OpenStudy (ranga):

You are just throwing out numbers.

OpenStudy (anonymous):

No, I'm not. I keep coming up with either 2036 or 2027 on my calculator.

OpenStudy (anonymous):

@PrincessLilian

OpenStudy (anonymous):

I'm not really an expert on this topic :/

OpenStudy (anonymous):

Okay, thanks anyways.

OpenStudy (anonymous):

I can try to help?

OpenStudy (anonymous):

Okay. That will work!

OpenStudy (anonymous):

This is so confusing I can't do it :( I think you should try subtracting the population by the percent. Then keep doing that for each year.

OpenStudy (anonymous):

Okay, thanks. Ill try that!

OpenStudy (anonymous):

:) I'm so sorry that I couldn't help.

OpenStudy (anonymous):

Its okay. No worries!

OpenStudy (ranga):

\(\Large t = \frac{\log(0.8539509)}{\log(0.989)} = ? \)

OpenStudy (anonymous):

I just had my mom try to do it. She got 2024. Im coming up with two different answers and she is coming up with another answer.

OpenStudy (ranga):

In the above calculation, t is approx. 14 and so 2010 + 14 = 2024 is correct.

OpenStudy (anonymous):

Okay great.